OpenStudy (anonymous):

x/5-18=-3 A. 3 B. 105 C. –105 D. 75

OpenStudy (anonymous):

add 18 to -3 x/5=15 now multiply 5 by 15

OpenStudy (anonymous):

x/ 5 - 18 = -3 x/ 5 - 18 + 18 = -3 + 18 x/ 5 = 15 x/ 5 * 5 = 15 * 5 x = 75

OpenStudy (anonymous):

Thanks Friend